David's Astronomy Pages
Notes - Session 1318 (2025-04-30)

 
Bullet Session Aims & Highlights
 - Observing Result
 - Night Summary Plot
 - Session Event Log
Bullet Operational Issues
  - Critical Issues (0),  Major Issues (0),  Minor Issues (0),  Small Defects (1),  Continuous Improvement (2)
Bullet Images from 2025-05-01  >>         
Bullet Geomagnetic Station - Remedial Work (2025-05-01 to 2025-05-07)
   
   

Session Aims & Highlights (2025-05-01)

Main aims

  1. Targets.  Acquire images of a selection of variable stars, nearby stars, comets & deep sky targets as allowed by sky conditions.
  2. Scope Path. Continue to test new Scope Path plotting routine and check that it is producing smoother/less angular paths

 Equipment & Software

Highlights

Notes

Summary Plots & Logs

Observing Plan
Image
  
Observing Result
Image
   
Image
 
Dome & Scope Slewing Performance
Image
  
Slew/Centering Performance
Image
  
Guiding Performance
Image
  
Sky Conditions (Locate Frames)
Image 
  
Night Sky Summary Plot
Top axis: Sky Brightness at Zenith (in ADU/s)
Lefthand axis: Local Time (hh LT). Righthand axis: Sun Altitude (degs)
Image   
  
Actual Weather vs Pre-Session Weather Forecast
Image
Image
  
Session Event Log
Time     Event Detail
22:22:31 Session Monitoring AutoStart monitoring for Live Session opportunity between 22:22 & 02:44
22:22:32   AutoStart Delayed Session delayed due to too few stars
22:27:32 Session AutoStarting Session autostarting (22:27)
22:27:53   CCDSoft Restarting CCDSoft being restarted (to set AutoSave No.)
22:28:20   Camera1 Connected SBIG Camera connected (Set point -15°C)
22:28:23 Session Created Session Created (Live, 2025-04-30 S01318, ImageSaveNum: 1318001)
22:28:28   Scope Switched On Telescope Power has been switched on via UPB Powerbox.
22:30:15   Services Started Observatory Services started
22:30:23 Observatory (Auto) Observatory placed in Fully-Automated Mode
22:30:25   AutoStart Delayed Session delayed due to too few stars
22:30:28 Session Pending Session pending (2025-04-30)
22:30:30 Session Initiating Session initiating (2025-04-30)
22:30:39   Camera1 Connected SBIG Camera connected (Set point -10°C)
22:30:46   Plan Requested Observing Plan requested from AstroPlan (1.43.4)
22:31:26   Plan Loaded Observing Plan loaded to queue (Plan ID: 1102)
22:32:07   Telescope Connected Telescope connected (TheSky6)
22:32:32 Session Equilibration Session ready to Open Dome
22:33:17   Dome Opened Dome opened (Opening time 45s, Zigbee 40s)
22:33:47 Session Running Session running
22:33:50   Queue Started Observing Queue started (18 targets selected)
22:33:52     Target Started (NrZen) Target started (Focus Field 12, HIP 58514)
22:34:06       Dome Unparked Dome unparked
22:38:00       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
22:41:03       Focusing Completed Foc1 AutoFocus Completed (Profile No 1, wide)
22:43:36       Focusing Completed Foc1 AutoFocus Completed (Profile No 1)
22:43:38       Focusing Started-Foc2 Foc2 Focusing Started (Secondary Scope, using ShCap)
22:43:54     Target Completed (NrZen) Target completed (Focus Field 12, HIP 58514)
22:47:14     Target Started (1/18) Target started (1/18, VV 410 w/SN2025hjz)
22:53:30       Focusing Skipped Foc1 focusing skipped - unable to find a star (TCF-S)
23:10:06     Target Completed Target completed (1/18, VV 410 w/SN2025hjz)
23:10:10     Target Started (2/18) Target started (2/18, AT2024accj)
23:15:00       Focusing Skipped Foc1 focusing skipped - star is too dim (TCF-S)
23:26:47       SoftSuspend Called Soft Suspend called due to Deteriorating Conditions (too few stars)
23:28:38       SoftSuspend Cancelled Soft Suspend is cancelled due to improved conditions
23:31:05     Target Completed Target completed (2/18, AT2024accj)
23:31:54     Target Started (3/18) Target started (3/18, GCVS RZ LMi)
23:35:48       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:38:24       Focusing Completed Foc1 AutoFocus Completed (Profile No 2)
23:43:49     Target Completed Target completed (3/18, GCVS RZ LMi)
23:43:54     Target Started (4/18) Target started (4/18, NGC 2407 w/SN2025gmc)
23:48:57       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:51:51       Focusing Completed Foc1 AutoFocus Completed (Profile No 3)
00:08:12     Target Completed Target completed (4/18, NGC 2407 w/SN2025gmc)
00:08:16     Target Started (5/18) Target started (5/18, 29P/Schwassmann-Wachmann)
00:11:44       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
00:14:25       Focusing Completed Foc1 AutoFocus Completed (Profile No 4)
00:29:53     Target Completed Target completed (5/18, 29P/Schwassmann-Wachmann)
00:29:57     Target Started (6/18) Target started (6/18, AT2025img (CV))
00:34:21       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
00:36:59       Focusing Completed Foc1 AutoFocus Completed (Profile No 5)
00:44:07     Target Completed Target completed (6/18, AT2025img (CV))
00:44:11     Target Started (7/18) Target started (7/18, UGC 5695 w/SN2025cfc)
01:05:10     Target Completed Target partially completed (7/18, UGC 5695 w/SN2025cfc)
01:05:14     Target Started (8/18) Target started (8/18, T CrB)
01:09:03       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:11:10       Focusing Completed Foc1 AutoFocus Completed (Profile No 6)
01:16:02     Target Completed Target completed (8/18, T CrB)
01:16:07     Target Started (9/18) Target started (9/18, HIP 87937)
01:18:28       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:21:09       Focusing Completed Foc1 AutoFocus Completed (Profile No 7)
01:24:53     Target Completed Target completed (9/18, HIP 87937)
01:24:57     Target Started (10/18) Target started (10/18, AT2025fma)
01:44:27     Target Completed Target completed (10/18, AT2025fma)
01:44:31     Target Started (11/18) Target started (11/18, 61 Cyg)
01:46:30       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:48:55       Focusing Completed Foc1 AutoFocus Completed (Profile No 8)
01:50:56     Target Completed Target completed (11/18, 61 Cyg)
01:51:00     Target Started (12/18) Target started (12/18, GCVS BL Lac)
01:54:12       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:56:47       Focusing Completed Foc1 AutoFocus Completed (Profile No 9)
02:01:15     Target Completed Target completed (12/18, GCVS BL Lac)
02:01:19     Target Started (13/18) Target started (13/18, WDS GRB 34)
02:05:30       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:08:03       Focusing Completed Foc1 AutoFocus Completed (Profile No 10)
02:10:24     Target Completed Target completed (13/18, WDS GRB 34)
02:10:28     Target Started (14/18) Target started (14/18, M31 w/Nova M31 2025-04c?)
02:13:45       Focusing Skipped Foc1 focusing skipped - star is lost (TCF-S)
02:30:14     Target Completed Target completed (14/18, M31 w/Nova M31 2025-04c?)
02:31:26     Target Started (15/18) Target started (15/18, UGC 12293 w/AT2025gyi)
02:35:44       Focusing Skipped Foc1 focusing skipped - star is too dim (TCF-S)
02:51:58   Critical Cloud Alert Critical Cloud Alert (Obs.Manager will close the Dome)
02:52:00   HardSuspend Called Hard Suspend called due to Critical Conditions (cloud)
02:52:03     Target Aborted (15/18) Target aborted (15/18, UGC 12293 w/AT2025gyi) due to cloud
02:52:05   Queue Paused Queue paused for Session Suspension
02:52:07 Session Suspending Session suspending
02:52:54   Dome Closed Dome closed (Closing time 50s, Zigbee 43s)
02:52:55 Session Suspended Session suspended
02:55:08 Session Resuming Session resuming
02:55:53     Target Started (NrZen) Target started (Focus Field 16, HIP 78946)
02:55:56   Dome Opened Dome opened (Opening time 45s, Zigbee 40s)
02:55:57 Session Running Session running
02:55:59   Queue Resumed Observing Queue resumed
03:00:14   Critical Cloud Alert Critical Cloud Alert (Obs.Manager will close the Dome)
03:00:16   HardSuspend Called Hard Suspend called due to Critical Conditions (cloud)
03:00:18     Target Aborted (NrZen) Target aborted (Focus Field 16, HIP 78946) due to cloud
03:00:20     Target Aborted (16/18) Target aborted (16/18, AT2024wqf) due to cloud
03:00:22   Queue Paused Queue paused for Session Suspension
03:00:24 Session Suspending Session suspending
03:01:09   Dome Closed Dome closed (Closing time 50s, Zigbee 43s)
03:01:12 Session Suspended Session suspended
03:13:02     Target Cancelled Target cancelled (16/18, AT2024wqf) due to cloud
03:24:14     Target Cancelled Target cancelled (17/18, GCVS HH And) due to cloud
03:43:11     Target Cancelled Target cancelled (18/18, GCVS EV Lac) due to cloud
03:43:15   Queue Completed Job Queue completed
03:43:19 Session Closing Session closing
03:43:23   Dome Closed Dome already closed (closed at 03:01, closing time 50s, zigbee 43s)
03:44:30   Dome Parked Dome parked (parking time 56s), Az: 90.0 deg
03:45:17   Telescope Parked Telescope parked (parking time 41s total)
03:45:40   Telescope State Handbox reads 'Scope parked Turn scope off.'
03:46:37   Telescope Switched Off Telescope Power has been switched off via UPB Switch.
03:48:01   Services Stopped Night Services stopped
03:48:04 Session Finishing Session Finishing started (Create Fits Summary, Transfer Files)
03:48:26     Dome (Find Park) Find Park started (Search Az 88.0 to 92.0°, Step 0.2°, Narrow)
03:49:10     Dome Find Park found (Best Park Az 90.4°)
03:49:16     Dome Dome has been synced at park position (Az 90.0°, Adjustment: -0.4°)
03:49:18   Dome Parked Dome parked
03:49:39   Dome Charging Dome is parked and re-charging Ok (13mA)
03:49:41 Session Finished Session Finished
 
Session Alerts & Alarms
Time     Type       Name Detail
03:49:39 Green Clear Dome Charging Dome is parked and re-charging Ok (13mA)
 

Back to Top


Operational Issues (2025-04-30, S1318)

[ Prev | Next ]

Critical Issues

Major Issues

Minor Issues

Small Defects

Continuous Improvement

[ Prev | Next ]

Back to Top


Geomagnetic Station -  Remedial Work (2025-05-01 to 2025-05-07)

A Geomagnetic Monitoring Station is being added to the Observatory in Spring 2025 for monitoring changes in the earth's magnetic field and to supplement and build upon the Observatory's existing Aurora monitoring that use a Black/White AllSky Camera and new NorthCam Colour Camera.

During Commissioning & Diagnositics phase it was discovered that the initial setup suffered from two significant issue (Noise & Anomalous Spiking) that require remediation.

The following notes describes the Remedial Work

Other notes  :
| Review | SAM III | Site Prep | Assembly Design | Parts Assembly | Testing | Final Assembly | Installation | Commissioning | Remedial Work |


Proposed Remedial Plan (2025-05-01) - For Discussion

The propsed remediation plan to resolve issues discovered during Commissioning Phase are outlined below

- dig up existing setup (downhole assembly and cables)
- retrieve existing magnetometers and temperature sensor
- build new assembly,
    - incorporate wider separation between magnetometers
    - connected using S/FTP shielded cable,
    - cable shielding will be grounded at sensor assemby end, by connecting it a common drain wire
       that will be run out of the assembly and embedded in the side of hole  (shielding won't be grounded at the Observatory end)
    - continue to use standard method using unbalanced circuits
- relocate the Temperature Sensor into its our buried enclosure at the same depth as the magnetometers, but separated by some distance.
- install new assembly
- connect cables back to SAM unit in Observatory

Alternative Remedial Plan - For Consideration

 -dig up existing cables from hole
- retrofit TIA-422 balanced circuit system above the
   - build and install outdoor interface box at sensor end incorporating TIA-422 balanced circuit board
   - build and install circuity at SAM III end using TIA-422 balanced circuit board
   - set jumper on SAM III to 12v
   - may need to lay & use use new cabling ? (how to ground if using shielded cabling)

potential issues
 - the time to extra complexity at both outdoor and indoor ends of the cable
-  the uncertainty on whether the spikes issue will be solved or not.
- uncertainty on making the outdoor enclosure be fully water proof..

Earlier Plan (from 2025-04-18) - On Hold

It seems that the Y Axis Magnetometer is compromised in some way (but needs to be confirmed with new tests on 2025-04-26).

Rather than dig up the existing 3-axis Sensor Fixture and pull out the working X & Z sensors and repair or install a new Y Axis magnetometer  it is proposed to build a dedicated new Y axis fixture for burial beside the existing fixture.

Requires
 - new FG-3+ magnetometer (£67.70 from https://robosavvy.co.uk/fgsensors-fg-3-sensor.html)
 - new fluxgate capacitors
      10 uF, 25v Tatalum Capacitor  (£ 1 + delivery)
     100 nF, 25v Ceramic Capacitor (£ 1 +delivery)
- new piping  (3 ends caps, 1 x 90 deg bend, 1 x 135 deg bend, 2 connectors - £10)
- new 15m length of Cat6 cable required (already available, but may opt for a shielded Cat6 cable given lessons / experience.

Work
  Assembly / test new magnetometer & fixture
  Part dig existing hole, install new Y Axis fixture
  Dig up trench,  thread new cablethrough corrugated pipe (if possible) or run cable beside corrugated piper


| Review | SAM III | Site Prep | Assembly Design | Parts Assembly | Testing | Final Assembly | Installation | Commissioning | Remediation |

Back to Top